Find the intergral curves of the equations
dx/yz=dy/zx=dz/xy
1
Expert's answer
2020-07-30T09:43:02-0400

Given equation is

"\frac{dx}{yz} = \frac{dy}{xz} = \frac{dz}{yx}"


We can write it as

"xdx = ydy = zdz"


Taking first two

"xdx = ydy"

Integrating both sides

"\frac{y^2}{2} = \frac{x^2}{2} + c_1 \implies y^2 - x^2 = C_1"


Taking last two

"ydy = zdz"

Integrating both sides

"\frac{z^2}{2} = \frac{y^2}{2} + c_2 \implies z^2 - y^2 = C_2"


Thus, "y^2-x^2=C_1" and "z^2-y^2=C_2" are integral curves.

Hence solution of the differential equation will be

"\Phi(y^2 - x^2, z^2 -y^2) = 0"

where "\Phi" is some arbitrary function.